# AceParse: A Comprehensive Dataset with Diverse Structured Texts for Academic Literature Parsing

### v1

@@ Abstract
Index Terms—  Academic literature parsing, Benchmark dataset, Vision-language model, Data-centric
1 Introduction
@@
The increasing focus on data-centric AI has shifted the emphasis from model development to the critical role of data quality in advancing AI technologies [1]. High-quality data are essential for training advanced models, particularly in domains requiring a deep understanding of complex information. Academic literature is a valuable data source due to its rich scientific content [2], but it is often stored in non-machine-readable formats like PDFs, necessitating effective parsing techniques to extract their information.
@@
Academic literature typically features a blend of structured content, including tables, formulas, lists, and algorithms, all of which work together to convey scientific insight. Parsing literature presents several key challenges. First, OCR-based methods [3, 4, 5] often focus on character recognition, leading to loss of structural information. Second, while modular approaches [6, 7, 8] can handle predefined content types like tables and formulas, they struggle with complex structures like algorithms and lists. Existing end-to-end parsing models, such as Nougat [9], are often trained on narrow proprietary datasets with a limited diversity of structured content, restricting their ability to generalize effectively across diverse structures. Finally, existing open-source datasets remain limited to character-level parsing [10, 11] or focus on specific content types like tables or formulas [12, 13], which does not cover the full diversity of structured elements present in academic documents.
@@
Fig. 1: (a) The construction process of the AceParse dataset consists of three stages: document collection, data synthesis, and boundary detection. (b) The network architecture of AceParser. Visual token embeddings and text token embeddings are concatenated into multimodal token embeddings, which are then processed by a BART-based multimodal encoder-decoder.
@@
To advance the unified parsing of academic literature containing various types of structured texts, we propose AceParse. Unlike previous datasets [10, 11] that either provide only character-level parsing results or focus on a single type of structured text [12, 13], AceParse encompasses a broad range of structured texts, including formulas, tables, lists, algorithms and sentences with embedded mathematical expressions. These texts are annotated using the LaTeX markup language to accurately describe their structure. As far as we know, AceParse is the first open-source dataset specifically designed for handling diverse structured content in academic literature parsing. A comparison of different datasets is provided in Table 1. Based on AceParse, we propose AceParser, a multimodal academic literature parsing model that adopts the network architecture referenced from Florence2 [14]. As shown in Table 2, our model is compared with other existing methods.
@@
2 METHODOLOGY
2.1 Dataset Construction
@@
Fig. 2: Statistics of the AceParse dataset. (a) The number of structured items comprising AceParse, including the number of structured sentences with embedded formulas, etc. (b) Frequency histogram indexed by the label character length. (c) Joint kernel density estimation of image dimensions.
Table 1: Comparison of different academic literature parsing datasets. ”Modality” indicates the type of content the dataset is designed for. ”Structured” refers to whether the labels are represented using markup languages. *AceParse includes various structured texts such as tables, formulas, lists, algorithms, etc.
@@
Table 2: Comparison of different academic literature parsing methods. ”Params” refers to the number of model parameters. ”Paradigm” describes whether the model uses a modular approach. ”Structured” indicates whether the method can parse structured texts.
@@
Fig. 3: Feature map and cross-attention matrices within the AceParser model. The matrices displayed on the right represent cross-attention from four layers of AceParser, with the top row showing attention before training and the bottom row after training. Structured text locations are highlighted with yellow ellipses.
Fig. 2 details the statistical attributes of the AceParse dataset. Each synthesized document image contains annotation text centered at 1,107 characters, matching the model’s processing capacity. As shown in Figure 2(c), the majority of images have dimensions centered at 974 × 493 pixels, with median values for width and height being 974 and 493 pixels, respectively. These dimensions ensure a balance between resolution and file size, optimizing both training efficiency and content clarity.
2.2 AceParser Network Architecture
@@
Table 3: Comparison of different parsing methods using various evaluation metrics. Metric abbreviations are used for clarity, such as LD for Levenshtein Distance.
@@
3 EXPERIMENTS AND DISCUSSIONS
3.1 Experimental Setup
@@
3.2 Comparison with Different Parsing Methods
As shown in Table 3, we compared various academic literature parsing methods on the AceParse dataset using evaluation metrics including Levenshtein Distance [18], BLEU [19], F1-score, and Jaccard Similarity [20]. Methods that are not structure-aware, such as Tesseract [3] and PPOCR [4], generally showed lower performance, due to their limited ability to handle markup languages. Among structure-aware methods, modular approaches like Pix2Text [6] and Mineru [7] were also less competitive, attributed to error accumulation across modules. Our end-to-end method achieved the best parsing results, with a 4.1-point improvement in the F1 score and a 5-point improvement in Jaccard similarity. However, a current limitation of our approach is its relatively slow parsing speed of 5.92 seconds per sample, which will be a focus for future optimization.
3.3 Case Study

### v2

@@ Abstract
Academic literature parsing, Benchmark dataset, Vision-language model, Data-centric
I Introduction
@@
The increasing focus on data-centric AI has shifted the emphasis from model to the critical role of data in advancing AI technologies [1, 2]. In this paradigm, the quality, diversity, and representativeness of data play a pivotal role in determining the success of AI systems. High-quality data are essential for training advanced models, particularly in domains requiring a deep understanding of complex information. Academic literature is a valuable data source due to its rich scientific content [3], but it is often stored in non-machine-readable formats like PDFs, necessitating effective parsing techniques to extract their information.
@@
Academic literature typically features a blend of structured content, including tables, formulas, lists, and algorithms, all of which work together to convey scientific insight. Parsing literature presents several key challenges. First, OCR-based methods [4, 5, 6] often focus on character recognition, leading to loss of structural information. These methods lack the ability to preserve the hierarchical and relational structure necessary for accurately understanding scientific content. Second, while modular approaches [7, 8] can handle predefined content types like tables and formulas, they struggle with complex structures like algorithms and lists. Moreover, the lack of integration between modules often results in inconsistent outputs for interconnected content types. Existing end-to-end parsing models, such as Nougat [9], are often trained on narrow proprietary datasets with a limited diversity of structured content, restricting their ability to generalize effectively across diverse structures. Finally, existing open-source datasets remain limited to character-level parsing [10, 11] or focus on specific content types like tables or formulas [12, 13], which does not cover the full diversity of structured elements present in academic documents. This gap highlights the pressing need for datasets that encompass a more comprehensive range of structured content to support advanced parsing techniques.
@@
Figure 1: (a) The construction process of the AceParse dataset consists of three stages: document collection, data synthesis, and boundary detection. (b) The network architecture of AceParser. Visual token embeddings and text token embeddings are concatenated into multimodal token embeddings, which are then processed by a BART-based multimodal encoder-decoder.
@@
To advance the unified parsing of academic literature containing various types of structured texts, we propose AceParse. Unlike previous datasets [10, 11] that either provide only character-level parsing results or focus on a single type of structured text [12, 13], AceParse encompasses a broad range of structured texts, including formulas, tables, lists, algorithms and sentences with embedded mathematical expressions. This comprehensive coverage addresses the limitations of existing datasets, enabling more versatile parsing solutions. Moreover, these texts are annotated using the LaTeX markup language to accurately describe their structure. As far as we know, AceParse is the first open-source dataset specifically designed for handling diverse structured content in academic literature parsing. A comparison of different datasets is provided in Table I. Based on AceParse, we propose AceParser, a multimodal academic literature parsing model that adopts the network architecture referenced from Florence2 [14]. As shown in Table II, our model is compared with other existing methods.
@@
II METHODOLOGY
II-A Dataset Construction
@@
Figure 2: Statistics of the AceParse dataset. (a) The number of structured items comprising AceParse, including the number of structured sentences with embedded formulas, etc. (b) Frequency histogram indexed by the label character length. (c) Joint kernel density estimation of image dimensions.
TABLE I: Comparison of different academic literature parsing datasets. ”Modality” indicates the type of content the dataset is designed for. ”Structured” refers to whether the labels are represented using markup languages. *AceParse includes various structured texts such as tables, formulas, lists, algorithms, etc.
@@
TABLE II: Comparison of different academic literature parsing methods. ”Params” refers to the number of model parameters. ”Paradigm” describes whether the model uses a modular approach. ”Structured” indicates whether the method can parse structured texts.
@@
Figure 3: Feature map and cross-attention matrices within the AceParser model. The matrices displayed on the right represent cross-attention from four layers of AceParser, with the top row showing attention before training and the bottom row after training. Structured text locations are highlighted with yellow ellipses.
Fig. 2 details the statistical attributes of the AceParse dataset. Each synthesized document image contains annotation text centered at 1,107 characters, matching the model’s processing capacity. As shown in Figure 2(c), the majority of images have dimensions centered at 974 $\times$ 493 pixels, with median values for width and height being 974 and 493 pixels, respectively. These dimensions ensure a balance between resolution and file size, optimizing both training efficiency and content clarity.
II-B AceParser Network Architecture
@@
TABLE III: Comparison of different parsing methods using various evaluation metrics. Metric abbreviations are used for clarity, such as LD for Levenshtein Distance.
@@
III EXPERIMENTS AND DISCUSSIONS
III-A Experimental Setup
@@
III-B Comparison with Different Parsing Methods
As shown in Table 3, we compared various academic literature parsing methods on the AceParse dataset using evaluation metrics including Levenshtein Distance [18], BLEU [19], F1-score, and Jaccard Similarity [20]. Methods that are not structure-aware, such as Tesseract [4] and PPOCR [5], generally showed lower performance, due to their limited ability to handle markup languages. Among structure-aware methods, modular approaches like Pix2Text and Mineru [7] were also less competitive, attributed to error accumulation across modules. Our end-to-end method achieved the best parsing results, with a 4.1-point improvement in the F1 score and a 5-point improvement in Jaccard similarity. However, a current limitation of our approach is its relatively slow parsing speed of 5.92 seconds per sample, which will be a focus for future optimization.
III-C Case Study